(4 - 3)[1 - (3 + 5)] x 5 = ?
Answer by stanbon(75887)

(4 - 3)[1 - (3 + 5)] x 5 = ?
-----------
1(1-8)*5 = ?
-7*5 = ?
? = -35
--------------
Cheers,
Stan H.
-----------